# Less is More: Pseudo-Label Filtering for Continual Test-Time Adaptation

### v1

@@ 4.2 Self-Adaptive Thresholding
$\tau_{t}(c)={\rm MaxNorm}(p_{t}(c))\cdot\tau_{t}=\frac{p_{t}(c)}{{\rm max}({p_{t}(c))}}\cdot\tau_{t},\quad t\neq 0,\forall c\in\mathcal{C}$ (3)
where MaxNorm is the Maximum Normalization (i.e., $x^{\prime}=\frac{x}{\max(x)}$), $\tau_{t}(c)$ represents the class-level adaptive threshold that combines the global threshold $\tau_{t}$ and class thresholds $p_{t}(c)$. Where the global threshold reflect overall model confidence and the class thresholds reflect class-specific confidence. In the following, we describe the design process of how to design initial threshold, global threshold, and class thresholds using three principles.
@@
$\tau_{t}=\left\{\begin{array}[]{ll}\vspace{1ex}{\rm EMA}(Q_{t},Q_{t-1})=\tau_{t-1}\cdot{\rm exp}(\alpha(\mathbb{E}_{\max}[{Q_{t-1}}]-\mathbb{E}_{\max}[{Q_{t}}])),&\mathbb{E}_{\max}[Q_{t}]<\mathbb{E}_{\max}[Q_{t-1}],\\ {\rm ED}(Q_{t})=\lambda\cdot\tau_{t-1}+(1-\lambda)\cdot\mathbb{E}_{\max}[{(Q_{t})}],&{\rm otherwise},\\ \end{array}\right.$ (5)
@@
$p_{t}(c)=\left\{\begin{array}[]{ll}\vspace{1ex}{\rm EMA}(Q_{t}(c),Q_{t-1}(c))&\mathbb{E}_{\max}[Q_{t}(c)]<\mathbb{E}_{\max}[Q_{t-1}(c)],\\ {\rm ED}(Q_{t}(c))&{\rm otherwise},\\ \end{array}\right.$ (6)
where $p_{t}(c)$ is the list containing all classes.
@@
$\mathcal{L}_{\rm u}=\frac{1}{B}\sum\nolimits_{b=1}^{B}{\mathbf{1}({\rm max}(Q)>\tau(c))}\cdot\mathcal{H}(q,Q),$ (7)
@@ 4.3 Class Prior Alignment
4:     Update the class threshold $p_{t}(c)$ (Eq. (6))
@@
$R_{t}=\frac{U(C)_{t}}{\mathbb{E}_{\max}[Q_{t}]}=\frac{{\rm Hist}_{B}[\mathbf{1}({\rm max}(q)>\tau(c))\cdot Q]}{\mathbb{E}_{\max}[{\mathbf{1}({{\rm max}(q)}>\tau(c))\cdot Q}]}.$ (8)
Similar to $R_{t}$, we compute $R_{s}$ as:
@@
$\tilde{h}=\lambda\tilde{h}+(1-\lambda){\rm Hist}_{B}(q_{s}).$ (10)
@@ 5.2 Main Results
For Testing On Principles. Take CIFAR10-C for example. Table.2 shows the advantage of low initial thresholds in the early target domain. Second, we test for different threshold and model confidence matching methods, including Fixed, DSS, Confidence Ratio Matching (CRM)+ED, and EMA+ED. Table.3 shows that positive correlation matching methods generally perform better. Finally, we compared the category granularity, as shown in Table.3, fine-grained thresholding can distinguish similar pseudo-labels more effectively.
@@ Appendix A Appendix / supplemental material
$P({Y}=1)=\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{2}}-{\mu_{1}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{2}}}})+\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{1}}-{\mu_{2}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{1}}}}),$
$P({Y}=-1)=\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{2}}-{\mu_{1}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{1}}}})+\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{1}}-{\mu_{2}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{2}}}}),$
@@ A.2 Derivation of Principle-2
To show that $P(Y=0)$ increases as $\tau$ gets bigger or $\beta$ gets smaller , we only need to show $P(Y=1||-1)$ gets smaller. First, we represent the irrelevant variables as constants, namely $a=\frac{\mu_{2}-\mu_{1}}{2},b_{1}=\frac{1}{2\sigma_{1}},b_{2}=\frac{1}{2\sigma_{2}},z=\frac{1}{\beta}$ Then, we denote $\log(\frac{\tau}{1-\tau})$ as $c$. Finally, due to the symmetry of $P(Y=1)$ and $P(Y=-1)${$b_{1}=b_{2}$} we obtain:
$P(Y=1)=P(Y=-1)=\frac{1}{2}\phi(ab-bcz)+\frac{1}{2}\Phi(-ab-bcz),$
@@
where $\phi$ is the probability density function of a standard normal distribution. Since$|ab-bcz|<|-ab-bcz|and|ab-bcz|<|-ab-bcz|$, we have $f^{\prime}(c)<0andh^{\prime}(z)<0$, and the proof-1 is complete.
@@ A.3 Derivation of Principle-3
Similarly, to show that $P(Y=0)$ increases as $z=\mu_{2}-\mu_{1}$ gets smaller, we only need to show $P(Y=1and-1)$ gets bigger. First, we represent the irrelevant variables as constants, namely $n_{1}=\frac{1}{2\sigma_{1}},n_{2}=\frac{1}{2\sigma_{2}},m_{2}=\frac{\log(\frac{\tau}{1-\tau})}{\sigma_{1}}$ Then, we denote $\mu_{2}-\mu_{1}$ as $z$. Finally, due to the symmetry of $P(Y=1)$ and $P(Y=-1)$($n=n_{1}=n_{2},m=m_{1}=m_{2}$)} we obtain:
$P(Y=1)=P(Y=-1)=\frac{1}{2}\Phi(nz-m)+\frac{1}{2}\Phi(-nz-m),$
@@ B.1 Filter ratio
$f(p)=\frac{\sum\nolimits_{n=1}^{N}{\mathbb{M}({\rm max}(q_{s})>\tau)}}{N}.$
@@ B.2 Quality
$g(p)=\frac{\mathbb{M}(({\rm max}(q_{s})>\tau)||\hat{p}=Y)}{\mathbb{M}(({\rm max}(q_{s})>\tau)}.$

### v2

@@ 4.2 Self-Adaptive Thresholding
$\tau_{t}^{*}(c)={\rm MaxNorm}(\tau_{t}(c))\cdot\tau_{t}=\frac{\tau_{t}(c)}{{\rm max}({\tau_{t}(c))}}\cdot\tau_{t},\quad t\neq 0,\forall c\in\mathcal{C}$ (3)
where MaxNorm is the Maximum Normalization (i.e., $x^{\prime}=\frac{x}{\max(x)}$), $\tau_{t}^{*}(c)$ represents the class-level adaptive threshold that combines the global threshold $\tau_{t}$ and class thresholds $\tau_{t}(c)$. Where the global threshold reflect overall model confidence and the class thresholds reflect class-specific confidence. In the following, we describe the design process of how to design initial threshold, global threshold, and class thresholds using three principles.
@@
$\tau_{t}=\left\{\begin{array}[]{ll}\vspace{1ex}{\rm EMA}(Q_{t})=\lambda\cdot\tau_{t-1}+(1-\lambda)\cdot\mathbb{E}_{\max}[{(Q_{t})}],&\mathbb{E}_{\max}[Q_{t}]>\mathbb{E}_{\max}[Q_{t-1}],\\ {\rm ED}(Q_{t},Q_{t-1})=\tau_{t-1}\cdot{\rm exp}(\alpha(\mathbb{E}_{\max}[{Q_{t-1}}]-\mathbb{E}_{\max}[{Q_{t}}])),&{\rm otherwise},\\ \end{array}\right.$ (5)
@@
$\tau_{t}(c)=\left\{\begin{array}[]{ll}\vspace{1ex}{\rm EMA}(Q_{t}(c))&\mathbb{E}_{\max}[Q_{t}(c)]>\mathbb{E}_{\max}[Q_{t-1}(c)],\\ {\rm ED}(Q_{t}(c),Q_{t-1}(c))&{\rm otherwise},\\ \end{array}\right.$ (6)
where $\tau_{t}(c)$ is the list containing all classes.
@@
$\mathcal{L}_{\rm u}=\frac{1}{B}\sum\nolimits_{b=1}^{B}{\mathbf{1}({\rm max}(Q)>\tau^{*}(c))}\cdot\mathcal{H}(q,Q),$ (7)
@@ 4.3 Class Prior Alignment
4:     Update the class threshold $\tau_{t}(c)$ (Eq. (6))
@@
$R_{t}=\frac{U(C)_{t}}{\mathbb{E}_{\max}[Q]}=\frac{{\rm Hist}_{B}[indi(Q)]}{\mathbb{E}_{\max}[indi(Q)]}.$ (8)
where $indi(\cdot)=\mathbf{1}({{\rm max}(\cdot)}>\tau^{*}(c))Q$ represent the probability of filter out pseudo-labels for the teacher model. Similar to $R_{t}$, we compute $R_{s}$ as:
@@
$\tilde{h}=\lambda\tilde{h}+(1-\lambda){\rm Hist}_{B}(q).$ (10)
@@ 5.2 Main Results
For Testing On Principles. Take CIFAR10-C for example. Table.2 shows the advantage of low initial thresholds in the early target domain. Second, we test for different threshold and model confidence matching methods, including Fixed, DSS, Confidence Ratio Matching (CRM) CRM+ED, and EMA+ED. Table.3 shows that positive correlation matching methods generally perform better. Finally, we compared the category granularity, as shown in Table.3, fine-grained thresholding can distinguish similar pseudo-labels more effectively.
@@ Appendix A Appendix / supplemental material
$P({Y_{p}}=1)=\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{2}}-{\mu_{1}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{2}}}})+\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{1}}-{\mu_{2}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{1}}}}),$
$P({Y_{p}}=-1)=\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{2}}-{\mu_{1}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{1}}}})+\frac{1}{2}\Phi(\frac{{\frac{{{\mu_{1}}-{\mu_{2}}}}{2}-\frac{1}{\beta}\log(\frac{{{\tau}}}{{1-{\tau}}})}}{{{\sigma_{2}}}}),$
@@ A.2 Derivation of Principle-2
To show that $P(Y_{p}=0)$ increases as $\tau$ gets bigger or $\beta$ gets smaller , we only need to show $P(Y_{p}=1||-1)$ gets smaller. First, we represent the irrelevant variables as constants, namely $a=\frac{\mu_{2}-\mu_{1}}{2},b_{1}=\frac{1}{2\sigma_{1}},b_{2}=\frac{1}{2\sigma_{2}},z=\frac{1}{\beta}$ Then, we denote $\log(\frac{\tau}{1-\tau})$ as $c$. Finally, due to the symmetry of $P(Y=1)$ and $P(Y_{p}=-1)${$b_{1}=b_{2}$} we obtain:
$P(Y_{p}=1)=P(Y_{p}=-1)=\frac{1}{2}\phi(ab-bcz)+\frac{1}{2}\Phi(-ab-bcz),$
@@
where $\phi$ is the probability density function of a standard normal distribution. Since$|ab-bcz|<|-ab-bcz|,|ab-bcz|<|-ab-bcz|$, we have $f^{\prime}(c)<0,h^{\prime}(z)<0$, and the proof-1 is complete.
@@ A.3 Derivation of Principle-3
Similarly, to show that $P(Y_{p}=0)$ increases as $z=\mu_{2}-\mu_{1}$ gets smaller, we only need to show $P(Y_{p}=1||-1)$ gets bigger. First, we represent the irrelevant variables as constants, namely $n_{1}=\frac{1}{2\sigma_{1}},n_{2}=\frac{1}{2\sigma_{2}},m_{2}=\frac{\log(\frac{\tau}{1-\tau})}{\sigma_{1}}$ Then, we denote $\mu_{2}-\mu_{1}$ as $z$. Finally, due to the symmetry of $P(Y_{p}=1)$ and $P(Y_{p}=-1)$($n=n_{1}=n_{2},m=m_{1}=m_{2}$)} we obtain:
$P(Y_{p}=1)=P(Y_{p}=-1)=\frac{1}{2}\Phi(nz-m)+\frac{1}{2}\Phi(-nz-m),$
@@ B.1 Filter ratio
$f(p)=\frac{\sum\nolimits_{n=1}^{N}{\mathbf{1}({\rm max}(q)>\tau)}}{N}.$
@@ B.2 Quality
$g(p)=\frac{\mathbf{1}(({\rm max}(q)>\tau)||\hat{p}=Y)}{\mathbf{1}(({\rm max}(q)>\tau)}.$